# EXCGEC: A Benchmark of Edit-wise Explainable Chinese Grammatical Error Correction

### v1

@@ Abstract
Existing studies explore the explainability of Grammatical Error Correction (GEC) in a limited scenario, where they ignore the interaction between corrections and explanations. To bridge the gap, this paper introduces the task of EXplainable GEC (EXGEC), which focuses on the integral role of both correction and explanation tasks. To facilitate the task, we propose EXCGEC, a tailored benchmark for Chinese EXGEC consisting of 8,216 explanation-augmented samples featuring the design of hybrid edit-wise explanations. We benchmark several series of LLMs in multiple settings, covering post-explaining and pre-explaining. To promote the development of the task, we introduce a comprehensive suite of automatic metrics and conduct human evaluation experiments to demonstrate the human consistency of the automatic metrics for free-text explanations. [^1] All the codes and data will be released after the review.
{CJK*}
UTF8gbsn
EXCGEC: A Benchmark of Edit-wise Explainable Chinese
@@
Jingheng Ye1thanks: indicates equal contribution., Shang Qin1∗, Yinghui Li1, Xuxin Cheng2, Libo Qin3 Hai-Tao Zheng1,4thanks: Corresponding author: Hai-Tao Zheng. (E-mail: zheng.haitao@sz.tsinghua.edu.cn), Peng Xing1, Zishan Xu1, Guo Cheng1, Zhao Wei5 1Tsinghua University, 2Peking University 3Central South University, 4Peng Cheng Laboratory, 5Tencent {yejh22,qin-s23}@mails.tsinghua.edu.cn
1 Introduction
Despite the notable advancements in Grammatical Error Correction (GEC) Bryant et al. (2023); Zhao et al. (2018); Ma et al. (2022); Ye et al. (2023a), there still exists a lack of profound examination into the explainability of GEC Dwivedi et al. (2023), which is critical in educational scenarios for L2 (Language second)-speakers Wang et al. (2021) or school-age children Li et al. (2023c). These mainstream users, who often face challenges in creating grammatically accurate and fluent texts, may be confused or even misguided if they are provided with limited access to only corrective texts. Therefore, augmenting the explainability of GEC is unquestionably beneficial for the progression of the GEC community as well as related fields, such as essay scoring Li et al. (2022b); Stahl et al. (2024), intelligent tutoring systems Montenegro-Rueda et al. (2023) and other emerging educational scenarios Li et al. (2024c); Du et al. (2024).
Figure 1: Task definitions of GEC, GEE and EXGEC. For the error description of EXGEC, we highlight evidence words, linguistic knowledge, error causes, and revision advice parts in different colors.
As illustrated in Figure 1, existing tasks like GEC and Grammatical Error Explanation (GEE) typically address either correction or explanation, ignoring the interaction between the two. To bridge the gap, we introduce the task of EXplainable Grammatical Error Correction (EXGEC). By integrating these two tasks, EXGEC enables systems to elucidate the linguistic knowledge and reasoning mechanism underlying predicted corrections, thereby achieving the best of both worlds. Additionally, EXGEC can function as a test bed for determining the explainable abilities of large language models (LLMs) and identifying any unintended biases and risks in educational scenarios.
To facilitate EXGEC, we present EXCGEC, a tailored benchmark for Chinese EXGEC, featuring the design of hybrid edit-wise explanations. Each explanation, based on a particular edit, consists of three elements: 1) Error types, which allow learners to absorb syntax and semantic knowledge in an inductive way Fei et al. (2023). We establish a hierarchical and pragmatic two-tier taxonomy for Chinese grammatical errors. 2) Error severity levels ranging from 1 $\sim$ 5 points, which are beneficial to prioritize core corrections. 3) Error descriptions, presented as the form of natural language explanation Camburu et al. (2018); He et al. (2023), provide evidence words, relevant linguistic knowledge or syntax rules, error causes, and revision advice for edits. The design provides more detailed and faithful guidance for learners, allowing them to comprehend each grammatical error committed. This is unlikely achievable for other designs such as example-based Kaneko et al. (2022) or sentence-level explanations Nagata et al. (2021).
Stimulated by recent success of synthetic data generation Shum et al. (2023); Whitehouse et al. (2023), we employ a semi-automatic dataset construction solution to enhance efficiency, while minimising annotation costs. Initially, we synthesize the EXCGEC dataset by prompting GPT-4 Liu et al. (2024). Then we hire native annotators to filter invalid data and provide a detailed analysis of invalid data, ensuring the quality of the dataset Ding et al. (2024). We finally obtain 8,216 clean explanation-augmented samples for benchmarking. We also introduce automatic metrics to evaluate performance across both tasks, and conduct human evaluation experiments to ascertain the correlation between these metrics and human judgement, thus demonstrating their effectiveness.
Based on the benchmark, we develop EXGEC baseline models that can perform both the correction and explanation tasks in either post-explaining (correct-then-explain) or pre-explaining (explain-then-correct) sequences. Particularly, we design Correct-Then-Explain (COTE) decoding algorithm for post-explaining models. Benchmarking various series of open-source LLMs has yielded several intriguing findings. For example, post-explaining models display higher performance than pre-explaining models. Moreover, COTE significantly enhances performance by alleviating the alignment workload for the LLMs. Our contributions in this paper are listed as follows:
  • (1)
    We introduce the EXGEC task and establish a corresponding benchmark consisting of a Chinese EXGEC dataset and a comprehensive set of automatic metrics, contributing to the stable development of the field of EXGEC.
  • (2)
  • (3)
2 Related Work
Exploration of explainable GEC has witnessed a paradigm shifting from fine-tuning Yang et al. (2020); Ye et al. (2022); Li et al. (2022c); Ye et al. (2023b); Kaneko and Okazaki (2023); Huang et al. (2023) to prompting Li et al. (2024b); Tan et al. (2023); Yu et al. (2024); Zhao et al. (2024); Xu et al. (2024); Qin et al. (2024a, b), with the focus being local explanations of individual predictions. Fei et al. (2023) construct an explainable GEC dataset EXPECT, which is annotated with evidence words and error types based on the standard GEC benchmark Bryant et al. (2019). However, EXPECT falls short of flexibility due to the lack of natural language explanations. To fill the gap, Song et al. (2023) propose the task of grammatical error explanation. They observe that GPT-4 suffers from identifying and explaining errors with limited access to only parallel source-target pairs. To address this issue, they fine-tune an extra LLM as an edit extractor, which is trained on synthesized data. However, all these studies overlook the benefits of effectiveness and efficiency brought by multi-task learning both correction and explanation tasks, which is extensively explored in this work.
On the other hand, a similar task, called the feedback comment generation task, focuses on sentence-level explanations Liu et al. (2022); Dong et al. (2023); Sun et al. (2022); Li et al. (2023b); Huang et al. (2024). However, it suffers from high cost associated with data annotation Nagata et al. (2020). Furthermore, due to the complexity of the task Nagata (2019), it is often explored with limited access to only a subset of grammatical error types.
3 Task Definition
3.1 Grammatical Error Correction
@@
GEC has been studied for decades, witnessing the shift from rule-based methods to LLM-based methods. Formally, given an ungrammatical text (source text) $X=\{x_{1},x_{2},\cdots,x_{T}\}$, a GEC model is required to correct $X$ into a grammatically correct counterpart (target text) $Y=\{y_{1},y_{2},\cdots,y_{T^{\prime}}\}$ without changing the original semantic as far as possible. Typically, GEC is usually treated as a sequence-to-sequence (Seq2Seq) task, the training objective of which is formulated as follows:
$\mathcal{L}_{\operatorname{GEC}}=-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X)$ (1)
3.2 Grammatical Error Explanation
@@
GEE has been explored in several methodologies, including sentence-level explanation and edit-wise explanation. Since sentence-level explanations suffer from over-generalization and confusion especially when a sentence contains multiple grammatical errors, this work focuses solely on edit-wise explanations. Given a source text $X$ and its target counterpart $Y$, the GEE model needs to explain each grammatical error $e_{i}$ in $X$. Specifically, GEE is typically solved in a two-step pipeline consisting of edit extraction and edit-wise explanation. 1) Edit extraction produces an edit set $E=\{e_{1},e_{2},\cdots,e_{n}\}$ that represent grammatical errors in $X$ and also clarify the transformation from ungrammatical segments of $X$ to target segments of $Y$. Typically, an edit contains four key elements: source position $sp$, source content $sc$, target position $tp$, and target content $tc$. The process of edit extraction can be easily accomplished using alignment-based evaluation toolkits like ERRANT Bryant et al. (2017); Felice et al. (2016) and CLEME Ye et al. (2023c). 2) Edit-wise explanation generates a set of explanations $E^{\prime}=\{e^{\prime}_{1},e^{\prime}_{2},\cdots,e^{\prime}_{n}\}$, with each explanation $e^{\prime}_{i}$ corresponding to $e_{i}$, given the source and the target texts. Although the design of explanation varies across related work Song et al. (2023); Zhao et al. (2024), the typical training objective of GEE models is presented as follows:
$E=f(X,Y)$ (2)
$\mathcal{L}_{\operatorname{GEE}}=-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i})$ (3)
@@
Existing studies Song et al. (2023); Fei et al. (2023) focus on developing GEE models that can generate more reasonable explanations. However, an extra GEC model is compulsory to allow GEE models to generate explanations if only source texts are offered, thus resulting in an issue of low efficiency. Furthermore, there exists a gap between GEC and GEE models if they are trained on different data with domain shifts.
Figure 2: Overview of benchmark construction and model development. We show the inference process of the post-explaining model in particular.
3.3 Explainable Grammatical Error Correction
To get rid of the drawbacks brought by the nature of GEE, we propose the EXGEC task which aims to perform both correction and explanation tasks simultaneously. The motivation for combining these two tasks majorly falls on two aspects. First, a branch of existing studies Wiegreffe and Marasovic (2021); Hartmann and Sonntag (2022); Li et al. (2022a, 2024a) have demonstrated training with access to human explanations can improve model performance. It is also intuitive that either of the GEC and GEE tasks can mutually benefit from each other when training in a multi-task manner. Second, it is more time-saving and cost-efficient to deploy a single EXGEC model rather than two detached models in foreign language education platforms.
@@ 3.3 Explainable Grammatical Error Correction
Models are trained first to generate target texts Camburu et al. (2018), which allows the explanations to be explicitly conditioned on the target texts, thus ensuring high faithfulness of explanations towards the target texts. The training objective is as follows:
$\mathcal{L}_{\operatorname{post}}=$ $-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X)$ (4)
$-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i})$
The inference of post-explaining models is represented as follows:
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{post}}(X)$ (5)
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{post}}(X,Y,f(X,\hat{Y}))$ (6)
With the target texts generated ahead, post-explaining models can output explanations conditioned on the specific edits that are extracted by aligning the source and the target texts, thus improving the accuracy and faithfulness of explanations.
@@
This type of models are trained conversely, whose mechanism is similar to the Chain of Thought (CoT) technique Chai et al. (2024). Pre-explaining models are supposed to make full use of synthesized explanations to generate elaborated target texts. With minimal modification from Equation (4), the training objective of pre-explaining models is as follows:
$\mathcal{L}_{\operatorname{pre}}=$ $-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X)$ (7)
$-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X,E^{\prime})$
@@
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{pre}}(X)$ (8)
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{pre}}(X,E^{\prime})$ (9)
4 EXCGEC Benchmark
To facilitate the development of EXGEC task, we construct EXCGEC, the first benchmark for explainable Chinese GEC particularly. As illustrated in Figure 2, we begin by the process of data curation, which consists of explanation design in Section 4.1, explanation synthesis and refinement in Section 4.2. Then we gain an in-depth understanding of GPT-4 Achiam et al. (2023) in EXGEC by further analyzing generated data in Section 4.3, where we summarize common failure modes in invalid instances. Finally, we introduce a series of automatic metrics for evaluating explanations in Section 4.4.
4.1 Explanation Design
In the pursuit of comprehensiveness and plausibility, we adopt a hybrid strategy for edit-wise explanations, where each edit is explained through three aspects, including error type labels, error severity levels, and free-text error descriptions. 1) Error type labels allow language learners to comprehend and infer syntax and grammar rules in an inductive manner. In particular, we employ a two-tier hierarchical taxonomy including 5 major types and 16 minor types shown in Table 1, inspired by existing studies Liping (2014); Peng et al. (2021); Zhang et al. (2022). The detailed descriptions of various error types are included in Appendix A.1 and A.2. If an edit covers multiple error types, we select the one with the highest granule. 2) Error severity levels, ranging from 1 to 5 points, indicate the significance of a specific grammatical error. 3) Error descriptions are the most crucial and flexible element. These provide keywords, pertinent linguistic knowledge, causes of errors, and revision guidance in a free-text format. We stipulate well-defined error descriptions should meet three principles: fluency, reasonability (making sense to humans), and faithfulness (targeted to a specific edit). To ensure reasonability and faithfulness, the error description must mostly conform to the syllogism form of deductive reasoning: [major premise: semantic rules and related knowledge], [minor premise: the reason for the error in the text], and [explain how to correct it]. Further, any evidence from the source $X$ must be enclosed within special markers 【 】. Similarly, correction content that occurs in the target sentence $Y$ must be enclosed within { }, as indicated in Figure 1.
@@ 4.1 Explanation Design
Spelling-level Error 字音混淆错误
Phonetic Confusion Error
字形混淆错误
Glyph Confusion Error
@@
Internal Character Misplacement Error
命名实体拼写错误
Named Entity Misspelling
@@
Table 1: Hierarchical taxonomy of grammatical error types defined in our benchmark.
4.2 Explanation Synthesizing
Annotating high-quality explanations on a large scale poses a huge challenge to our benchmark construction. Hence, we leverage GPT-4 to synthesize edit-wise explanations efficiently. To achieve this, we first select 10,000 parallel samples across 6 existing benchmarks or datasets of Chinese GEC, with the details listed in Table 2. We pick out only the samples with changed target sentences and select the single target sentence with the most edits if a sample is annotated with multiple target sentences. Then, we prompt GPT-4 to generate edit-wise explanations following in-context learning. To ensure the faithfulness of the synthesized explanation, we first extract edits using the toolkit CLEME Ye et al. (2023c). Inspired by Li et al. (2022a), we then employ the Rationalization Prompting (RP) strategy, where we concatenate task definition, demonstrations, and a parallel sample $(X,Y)$ with extracted edits $E=\{e_{1},e_{2},\cdots,e_{n}\}$ as the prompt. For each error type, we provide the definition, a suggested template of error description, and a demonstration. The prompt is listed in Appendix A.3.
@@ 4.2 Explanation Synthesizing
4.3 Explanation Refinement and Analysis
Benefiting from the extensive knowledge acquired during the large-scale pre-training process, GPT-4 is able to generate fluent, reasonable, and plausible explanations in most cases, meeting the requirements with specified instructions. However, GPT-4 is not guaranteed to produce all high-quality explanations due to hallucination, and the patterns of those invalid explanations are referred to as failure modes. Therefore, we hire 12 native speakers, all of whom are graduated students, to screen out invalid explanations. We finally obtain 8,216 clean samples out of 10,000 samples. We further investigate the failure modes of invalid explanations generated by GPT-4, which is provided in Appendix A.4.
4.4 Automatic Metrics
Recent studies leverage human evaluation for the evaluation of GEE due to the lack of enough annotated samples, posing a challenge for the efficient development of EXGEC systems. In this paper, we introduce a comprehensive set of automatic metrics for both correction and explanation parts.
@@ 4.4 Automatic Metrics
We employ CLEME and ChERRANT to evaluate the correction performance. Both are edit-based metrics that output P/R/F0.5 scores, and they have been proven reliable metrics for GEC on CoNLL-2014 Ye et al. (2023c).
@@
Since an edit-wise explanation consists of three critical elements, we define respectively automatic metrics for them. 1) Accuracy and Macro-F1 scores are computed for error type clarification, following the conventional evaluation protocol of text clarification Li et al. (2020). 2) We report the mean absolute error (MAE) to show the deviation of hypothesis error severity levels towards ground truth ones. 3) We employ various metrics for evaluating the free-text explanation description, including BLEU, METEOR, ROUGE-1, ROUGE-2, and ROUGE-L. We leave the analysis on the effectiveness of these metrics to Section 7.2.
5 Method
5.1 Training
To streamline the training process covering all the tasks mentioned in Section 3, we treat all of them as a unified Seq2Seq task Yang et al. (2022b, a). To achieve this, we linearize the data in the format of json Gao et al. (2023). This structured approach simplifies the process of output parsing involving three types elements of edit-wise explanations, and provides a consistent and controllable view to distinguish tasks, enabling the model to understand essential task elements and their relations. Therefore, we train all models using the same smooth cross-entropy loss, regardless of the specific task.
5.2 Inference

### v2

# EXCGEC: A Benchmark for Edit-Wise Explainable Chinese Grammatical Error Correction

@@ Abstract
Existing studies explore the explainability of Grammatical Error Correction (GEC) in a limited scenario, where they ignore the interaction between corrections and explanations and have not established a corresponding comprehensive benchmark. To bridge the gap, this paper first introduces the task of EXplainable GEC (EXGEC), which focuses on the integral role of correction and explanation tasks. To facilitate the task, we propose EXCGEC, a tailored benchmark for Chinese EXGEC consisting of 8,216 explanation-augmented samples featuring the design of hybrid edit-wise explanations. We then benchmark several series of LLMs in multi-task learning settings, including post-explaining and pre-explaining. To promote the development of the task, we also build a comprehensive evaluation suite by leveraging existing automatic metrics and conducting human evaluation experiments to demonstrate the human consistency of the automatic metrics for free-text explanations. Our experiments reveal the effectiveness of evaluating free-text explanations using traditional metrics like METEOR and ROUGE, and the inferior performance of multi-task models compared to the pipeline solution, indicating its challenges to establish positive effects in learning both tasks.
Code & Datahttps://github.com/THUKElab/EXCGEC
Introduction
Despite the notable advancements in Grammatical Error Correction (GEC) (Bryant et al. 2023; Ye et al. 2023a; Li et al. 2025), there still exists a lack of profound examination into the explainability of GEC (Dwivedi et al. 2023), which is critical in educational scenarios for L2 (second language)-speakers (Wang et al. 2021). These mainstream users, who often face challenges in creating grammatically accurate and fluent texts, may be confused or even misguided if provided with limited access to only corrective texts (Ye et al. 2025). Therefore, augmenting the explainability of GEC is unquestionably beneficial for the progression of GEC as well as related fields, such as essay scoring (Stahl et al. 2024), intelligent tutoring systems (Montenegro-Rueda et al. 2023).
Figure 1: Task definitions of GEC, GEE, and EXGEC. We highlight 【evidence words】, {correction}, linguistic knowledge, error causes, and revision advice parts.
As illustrated in Figure 1, existing tasks like GEC (Ye et al. 2022) and Grammatical Error Explanation (GEE) typically address either correction or explanation, ignoring the interaction between the two. To bridge the gap, we introduce the task of EXplainable Grammatical Error Correction (EXGEC). By integrating these two tasks, EXGEC enables systems to elucidate the linguistic knowledge and reasoning mechanism underlying predicted corrections, thus achieving the best of both worlds. Additionally, EXGEC can function as a test bed for determining the explainable abilities of large language models (LLMs) and identifying any unintended biases and risks in educational scenarios.
To facilitate EXGEC, we present EXCGEC, a tailored benchmark for Chinese EXGEC, featuring the design of hybrid edit-wise explanations. Each explanation, based on a particular edit, consists of three elements: 1) Error types, which allow learners to absorb syntax and semantic knowledge in an inductive way (Fei et al. 2023). We establish a hierarchical and pragmatic two-tier taxonomy for Chinese grammatical errors. 2) Error severity levels ranging from 1 $\sim$ 5 points, which are beneficial to prioritize core corrections. 3) Error descriptions, presented as the form of natural language explanation (Camburu et al. 2018; He et al. 2023), provide evidence words, relevant linguistic knowledge or syntax rules, error causes, and revision advice for edits. The edit-wise design provides more detailed and faithful guidance for learners, allowing them to comprehend each grammatical error committed. This is unlikely achievable for other designs such as example-based (Kaneko et al. 2022) or sentence-level explanations (Nagata et al. 2021).
Stimulated by the recent success of synthetic data generation (Shum, Diao, and Zhang 2023; Whitehouse, Choudhury, and Aji 2023), we employ a semi-automatic dataset construction solution to enhance efficiency, while minimising annotation costs. Initially, we synthesize the evaluation part of EXCGEC by prompting GPT-4 (Liu et al. 2024). Then we hire experienced annotators to filter out invalid data and concurrently provide a detailed analysis of the invalid data, ensuring the quality of our dataset (Ding et al. 2024). We finally obtain 8,216 clean explanation-augmented samples for benchmarking. Additionally, We utilize existing automatic metrics to evaluate the performance. Particularly for error descriptions, we conduct a human evaluation experiment to ascertain the correlation between the metrics and human judgements, thus demonstrating their effectiveness.
Based on the benchmark, we develop EXGEC multi-task baseline models that can perform both the correction and explanation tasks in either post-explaining (correct-then-explain) or pre-explaining (explain-then-correct) sequences. Particularly, we design Correct-Then-Explain (COTE) decoding algorithm for post-explaining models. Benchmarking various series of open-source LLMs has yielded several intriguing findings. For example, post-explaining models display higher performance than pre-explaining models. However, both of them under-perform the pipeline solution. Moreover, COTE significantly enhances performance by alleviating the alignment workload for the LLMs. Our contributions in this paper are listed as follows:
  • We introduce the EXGEC task and establish a corresponding benchmark consisting of a Chinese EXGEC dataset and a comprehensive set of metrics, contributing to the stable development of the field of EXGEC.
Related Work
Explainable GEC.
Exploration of explainable GEC has witnessed a paradigm shift from fine-tuning to prompting (Zhao et al. 2024). EXPECT (Fei et al. 2023) is an explainable GEC dataset annotated with evidence words and error types based on the standard GEC benchmark (Bryant et al. 2019). However, EXPECT falls short of flexibility due to the lack of natural language explanations. To fill the gap, Song et al. (2023) propose the task of grammatical error explanation. They observe that GPT-4 suffers from identifying and explaining errors with limited access to only parallel source-target pairs. To address this issue, they fine-tune an extra LLM as an edit extractor trained on synthesized data. On the other hand, a similar task called feedback comment generation, focuses on sentence-level explanations. However, it suffers from expensive costs associated with data annotation (Nagata, Inui, and Ishikawa 2020). Furthermore, it is explored with limited access to only a subset of English grammatical error types due to the complexity of the task (Nagata 2019). In conclusion, all these studies do not establish a comprehensive benchmark integrating both the tasks of GEC and GEE, and thus lack in-depth exploration in multi-task learning the both tasks. However, our work is the first to propose a systematic framework for EXCGEC.
Chinese GEC.
The research on CGEC (Ye et al. 2023a; Ye, Li, and Zheng 2023) has also come a long way recently, along with a series of CGEC datasets (Zhao et al. 2018). Similar to those in English, Chinese grammatical errors can also be categorized into different error types. CLG (Ma et al. 2022) divides Chinese grammatical errors into 6 categories: Structural Confusion, Improper Logicality, Missing Component, Redundant Component, Improper Collocation, and Improper Word Order. However, the taxonomy of CLG is targeted toward grammatical errors made by native speakers and thereby can not cover those made by L2 speakers. To fill the gap, we design a two-tier hierarchical taxonomy, which is capable of covering most grammatical errors.
Task Definition
@@ Introduction
GEC (Schneider and McCoy 1998) has been studied for decades, witnessing the shift from rule-based methods to LLM-based methods. Formally, given an ungrammatical text (source text) $X=\{x_{1},x_{2},\cdots,x_{T}\}$, a GEC model is required to correct $X$ into a grammatically correct counterpart (target text) $Y=\{y_{1},y_{2},\cdots,y_{T^{\prime}}\}$ without changing the original semantic as far as possible. Typically, GEC is usually treated as a sequence-to-sequence (Seq2Seq) task, the training objective of which is formulated as follows:
$\mathcal{L}_{\operatorname{GEC}}=-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X).$ (1)
Grammatical Error Explanation
@@
GEE (Song et al. 2023) has received much attention recently and has been explored in several methodologies, including sentence-level explanation and edit-wise explanation. Since sentence-level explanations suffer from over-generalization and confusion especially when a sentence contains multiple grammatical errors, this work focuses solely on edit-wise explanations. Given a source text $X$ and its target counterpart $Y$, the GEE model needs to explain each grammatical error $e_{i}$ in $X$. Specifically, GEE is typically solved in a two-step pipeline consisting of edit extraction and edit-wise explanation. 1) Edit extraction produces an edit set $E=\{e_{1},e_{2},\cdots,e_{n}\}$ that represent grammatical errors in $X$ and also clarify the transformation from ungrammatical segments of $X$ to target segments of $Y$. Typically, an edit contains four key elements: source position $sp$, source content $sc$, target position $tp$, and target content $tc$. The process of edit extraction can be easily accomplished using alignment-based evaluation toolkits like ERRANT (Bryant, Felice, and Briscoe 2017) and CLEME (Ye et al. 2023b, 2024). 2) Edit-wise explanation generates a set of explanations $E^{\prime}=\{e^{\prime}_{1},e^{\prime}_{2},\cdots,e^{\prime}_{n}\}$, with each explanation $e^{\prime}_{i}$ corresponding to $e_{i}$, given $X$ and $Y$. Although the design of explanation varies across related work (Song et al. 2023; Zhao et al. 2024), the typical training objective of GEE models is presented as follows:
$E=f(X,Y),$ (2)
$\mathcal{L}_{\operatorname{GEE}}=-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i}),$ (3)
@@
Existing studies (Song et al. 2023; Fei et al. 2023) focus on developing GEE models that can generate explanations. However, an extra GEC model is compulsory for GEE models to work, thus resulting in an issue of low efficiency.
Figure 2: Overview of the benchmark and the model. We show the inference process of a post-explaining model in particular.
Explainable Grammatical Error Correction
To get rid of the drawbacks brought by the nature of GEE, we propose the EXGEC task which aims to perform both correction and explanation tasks simultaneously. The motivation for combining these two tasks majorly falls on two aspects. First, a branch of existing studies (Wiegreffe and Marasovic 2021; Hartmann and Sonntag 2022; Li et al. 2022, 2024) have demonstrated training with access to human explanations can improve model performance. It is also intuitive that either of the GEC and GEE tasks can mutually benefit from each other when training in a multi-task manner. Second, it is more time-saving and cost-efficient to deploy a single EXGEC model rather than two detached models in foreign language education platforms.
@@
Models are trained first to generate target texts (Camburu et al. 2018), which allows the explanations to be explicitly conditioned on the target texts, thus ensuring high faithfulness of explanations towards the target texts. The training objective is as follows:
$\mathcal{L}_{\operatorname{post}}=$ $-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X)-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X,Y,e_{i}).$ (4)
The inference of post-explaining models is as follows:
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{post}}(X),$ (5)
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{post}}(X,Y,f(X,\hat{Y})).$ (6)
With the target texts generated ahead, post-explaining models can output explanations conditioned on the specific edits extracted by an aligning process, thus improving the accuracy and faithfulness of explanations.
@@
This type of model is trained in converse order, whose mechanism is similar to the Chain of Thought (CoT) technique. Pre-explaining models are supposed to make full use of synthesized explanations to generate elaborated target texts. With minimal modification from Equation (4), the training objective of pre-explaining models is as follows:
$\mathcal{L}_{\operatorname{pre}}=$ $-\sum_{i=1}^{n}\log P(e^{\prime}_{i}\mid X)-\sum_{t=1}^{T^{\prime}}\log P(y_{t}\mid Y_{<t},X,E^{\prime}).$ (7)
@@
$\hat{E}^{\prime}=\operatorname{EXGEC}_{\operatorname{pre}}(X),$ (8)
$\hat{Y}=\operatorname{EXGEC}_{\operatorname{pre}}(X,E^{\prime}).$ (9)
EXCGEC Benchmark
To facilitate the development of EXGEC task, we construct EXCGEC, the first benchmark for explainable Chinese GEC particularly. As illustrated in Figure 2, we begin with the process of data curation, which consists of Explanation Design, Explanation Synthesizing, Explanation Refinement, and Analysis. Then we gain an in-depth understanding of GPT-4 (Achiam et al. 2023) by further analyzing the generated explanations, where we summarize common failure modes in invalid instances. Finally, we explain the evaluation for both the correction and the explanation tasks.
@@
Spelling-level Error 字音混淆错误 (Phonetic Confusion Error)
字形混淆错误 (Glyph Confusion Error)
@@
(Internal Character Misplacement Error)
命名实体拼写错误 (Named Entity Misspelling)
@@
Table 1: Hierarchical taxonomy of grammatical error types.
Explanation Design
In the pursuit of comprehensiveness and plausibility, we adopt a hybrid strategy for edit-wise explanations, where each edit is explained through three aspects, including error type labels, error severity levels, and free-text error descriptions. 1) Error type labels allow language learners to comprehend and inductively infer syntax and grammar rules. In particular, we employ a two-tier hierarchical taxonomy including 5 major types and 16 minor types shown in Table 1, inspired by authoritative linguistic books (Huang and Liao 2011; Shao 2016). Detailed descriptions of various error types are included in the supplementary materials. If an edit covers multiple error types, we select the one with the highest granule. 2) Error severity levels, ranging from 1 to 5 points, indicate the significance of a specific grammatical error. 3) Error descriptions are the most crucial and flexible element. These provide keywords, pertinent linguistic knowledge, causes of errors, and revision guidance in a free-text format. We stipulate well-defined error descriptions should meet three nonoverlapping principles: fluency, reasonability (making sense to humans), and faithfulness (targeted to a specific edit). To ensure reasonability and faithfulness, the error description must mostly conform to the syllogism form of deductive reasoning: [major premise: semantic rules and related knowledge], [minor premise: the reason for the error in the text], and [explain how to correct it]. Further, any evidence from the source $X$ must be enclosed within special markers 【 】. Similarly, correction content that occurs in the target sentence $Y$ must be enclosed within { }, as indicated in Figure 1.
Explanation Synthesizing
Annotating high-quality explanations on a large scale poses a huge challenge to our benchmark construction. Hence, we leverage GPT-4 to synthesize edit-wise explanations efficiently. To achieve this, we first select 10,000 parallel samples across 6 existing benchmarks or datasets of Chinese GEC, including FCGEC (Xu et al. 2022), YACLC (Wang et al. 2021), MuCGEC (Zhang et al. 2022), NaCGEC (Ma et al. 2022), NLPCC (Zhao et al. 2018) and HSK (Zhang 2009). The details are listed in Table 2. We pick out only the samples with changed reference sentences to maximize training efficiency (Zhang et al. 2022). We select the reference sentence with the most edits as the target sentence if a sample is annotated with multiple reference sentences. Then, we prompt GPT-4 to generate edit-wise explanations following in-context learning. To ensure the faithfulness of the synthesized explanation, we first extract edits using the toolkit CLEME (Ye et al. 2023b). Inspired by Li et al. (2022), we then employ the Rationalization Prompting (RP) strategy, where we concatenate task definition, demonstrations, and a parallel sample $(X,Y)$ with extracted edits $E=\{e_{1},e_{2},\cdots,e_{n}\}$ as the prompt. For each error type, we provide the definition, a suggested template of error description, and a demonstration. The prompt is listed in the supplementary materials.
@@
Explanation Refinement and Analysis
Benefiting from the extensive knowledge acquired during the large-scale pre-training process, GPT-4 can generate fluent, reasonable, and plausible explanations in most cases, meeting the requirements with specified instructions. However, GPT-4 is not guaranteed to produce all high-quality explanations due to hallucination, and the patterns of those invalid explanations are referred to as failure modes. Therefore, we hired 12 native speakers, all of whom are Chinese post-graduated students specializing in Chinese linguistics, to screen out invalid explanations. Before formal annotation, we compile the annotation guidelines and all the annotators receive intensive training. Two authors of the paper, who are also in charge of compiling the annotation guidelines, have made sure that their annotation accuracies are over 90% on testing samples. We make sure that each formal sample is checked by at least two annotators. We finally obtained 8,216 clean samples out of 10,000 samples. We further investigate the failure modes of these invalid explanations, which are provided in the supplementary materials.
Automatic Metrics
To promote the efficient development of EXGEC systems, we introduce a comprehensive suite of automatic metrics for both correction and explanation parts. Additionally, we conduct a human evaluation experiment in Section Analysis to demonstrate the alignment of the metrics used for assessing error descriptions with human judgments.
@@
We employ CLEME (Ye et al. 2023b) and ChERRANT (Zhang et al. 2022) to evaluate the correction performance. Both are edit-based metrics that output P/R/F0.5 scores, which have been proven reliable metrics for GEC on CoNLL-2014 (Ye et al. 2023b).
@@
Since an edit-wise explanation consists of three critical elements, we define respectively automatic metrics for them. 1) Accuracy and Macro-F1 scores are computed for error type clarification, following the conventional evaluation protocol of text clarification (Li et al. 2020). 2) We report the mean absolute error (MAE) to show the deviation of hypothesis error severity levels towards ground truth ones. 3) We employ various metrics for evaluating the free-text explanation descriptions considering both the reproductivity and efficiency, including BLEU (Papineni et al. 2002; Clinciu, Eshghi, and Hastie 2021), METEOR (Banerjee and Lavie 2005), ROUGE (Lin 2004).
Method
Training.
To streamline the training process covering all the tasks mentioned previously, we treat all of them as a unified Seq2Seq task. To achieve this, we linearize the data in the format of JSON (Gao et al. 2023). This structured approach simplifies the process of output parsing involving three elements of edit-wise explanations, and provides a consistent and controllable view to distinguish tasks, enabling the model to understand essential task elements and their relations. With this uniform format stipulation, we can train all models using the same smooth cross-entropy loss, regardless of the specific task.